Meeting 2022 01 05
From Hall A Wiki
GMn GEM commissioning Meetings: WEDNESDAY 10:00am - 11:00am
Link to all previous meetings: GMn GEM Commissioning Meeting
Please use sbs_gems mailing list for plots/communications and HALOG entries relevant to GEMs!
Update on BB GEMs(Sean, Ezekiel, Holly, Anu)
- Electronics swapped from BB layers on Monday.
- Power outage yesterday in EEL and test lab.
- Moved new UV layer over to test lab this morning.
- Work plan today to take pedestals followed by cosmics in test lab overnight. Also debugging a few APVs in the Hall.
- Could still do the swap tomorrow, with Friday as a backup. Probable work over the weekend to check everything.
- Still working on a long HV cable for the Hall. Mark Taylor to solder resistor on new UV layer today.
Update on UVa GEM layers for GEn-RP (Anu, John, Thir)
Update on INFN GEM layers for GEn-RP (Ezekiel)
- Layers ready for more cosmic testing.
- Still need to resolve and HV-shielding issue on J1 top.
Update on the Tracking efficiency / resolution (Andrew)
- Farm jobs are backed up.
Update on the High rate DAQ development (Ben, Bryan, Alex, Holly)
- Request to be able to run DAQ even when APVs are missing.